OpenStudy (anonymous):

solve the system of equations 2x-2y+z=-7 -x+3y+2z=0 x-4y-3z=1

OpenStudy (abb0t):

How about you solve this first: \(\left\{\begin{matrix} \frac{1}{x^4}+\frac{1}{y^4} =6& & \\ \frac{2}{x^4}-\frac{2}{y^4}=19& & \end{matrix}\right.\)

OpenStudy (anonymous):

._. what is that

OpenStudy (abb0t):

system of equations.

OpenStudy (anonymous):

I don't understand your methods.

OpenStudy (mathmate):

@abb0t Does your problem have anything to do with needy's problem? If not, it is recommended that you make a new post.

OpenStudy (mathmate):

@needy>.< You will have to solve the system of equations by either elimination, or by reduction. Do you know any of these methods?

OpenStudy (anonymous):

elimination yes